xx HydroInitExcision - Initial Excision for Hydro Evolutions

HydroxInitExcision - Initial Excision for Hydro Evolutions

Andrea Nerozzi, Frank Lxffler

May 16, 2024

Abstract

This thorn initialises the excision-spacemask for GRHydro to specified forms. It is based on a similar version used to initialise the excision-mask for Whisky code.

1 Using This Thorn Using this thorn is quite easy: You have to activate it in your parameter file and specify at least the following:

hydro_initexcision::hydro_initexcisionx=x"yes"

In addition to that you may want to specify the type of excision zone you want to have. You can do this by

hydro_initexcision::hydro_initexcision_typex=x<TYPE>

where

<TYPE>

is one of the following list: ”x-axis” ”y-axis” ”z-axis” ”box” ”sphere”

2 Thorn Source Code This was initially written by Andrea Nerozzi and later changed by Frank Lxffler.

3 Thorn Documentation This documentation was largely written by Frank Lxffler.

4 Parameters

11hydroxinitexcision 10Scope: restrictedx01 BOOLEAN

11Description: Sets up an excision zone

11                             10  x01  Default: no

11

11hydroxinitexcisionxcoordinatexlength 10Scope: restrictedx01  REAL

11Description: Length of excised region

11 Range 10 x01 Default: (none)
110:  10Length of excised region, has different meanings for the types

11

11hydroxinitexcisionxflip  10Scope: restrictedx01  INT

11Description: Exchange excised and normal region

11 Range 10 x01 Default: (none)
110:1  100 (default) for no, 1 for yes

11

11hydroxinitexcisionxfraction  10Scope: restrictedx01  REAL

11Description: Fraction of excised region

11 Range 10 x01 Default: 0.25
110:1  10Represents the fraction of excised region (0..1)

11

11hydroxinitexcisionxminxpoints 10Scope: restrictedx01  INT

11Description: minimal number of points of excision

11 Range 10 x01 Default: (none)
110:  10minimal number of points of excision, currently only partly implemented

11

11hydroxinitexcisionxoldxmask  10Scope: restrictedx01  INT

11Description: Also set the old ’emask’?

11 Range 10 x01 Default: (none)
110:1  100 (default) for no, 1 for yes

11

11hydroxinitexcisionxpositionxx 10Scope: restrictedx01  REAL

11Description: x-coordinate

11 Range 10 x01 Default: 0.0
11:  10anything real is ok

11

11hydroxinitexcisionxpositionxy 10Scope: restrictedx01  REAL

11Description: y-coordinate

11 Range 10 x01 Default: 0.0
11:  10anything real is ok

11

11hydroxinitexcisionxpositionxz 10Scope: restrictedx01  REAL

11Description: z-coordinate

11 Range 10 x01 Default: 0.0
11:  10anything real is ok

11

11hydroxinitexcisionxpoststep  10Scope: restrictedx01  BOOLEAN

11Description: Schedule also at poststep

11                             10  x01  Default: no

11hydroxinitexcisionxtype  10Scope: restrictedx01  KEYWORD

11Description: Type of excision to apply

11 Range 10 x01 Default: x-axis
11box  10Box excised region
11x-axis  10X axis excision
11y-axis  10Y axis excision
11z-axis  10Z axis excision
11diagonal  10diagonal excision
11sphere  10sphere excision

11

11domain  10Scope: shared from GRIDx01KEYWORD

11

5 Interfaces

General Implements: hydroxinitexcision

Inherits: hydrobase grid spacemask

6 Schedule This section lists all the variables which are assigned storage by thorn EinsteinInitialData/HydroxInitExcision. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxINITIAL x hydroxinitexcisionmask x sets up the excision mask x

11x11After: 11maskone
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxPOSTxRECOVERxVARIABLES x hydroxinitexcisionmask x sets up the excision mask x

11x11After: 11maskone
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxPOSTREGRID x hydroxinitexcisionmask x sets up the excision mask x

11x11After: 11maskone
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxPOSTSTEP x hydroxinitexcisionmaskxpoststep x sets up the excision mask x

11x11Before: 11setupxepsdis
11 x 11Language:11c
11 x 11Type: 11function